# Raoyuan Zhao MaiNLP, Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML)

### v2

@@ Abstract
Large language models (LLMs) are used globally across many languages, but their English-centric pretraining raises concerns about cross-lingual disparities for cultural awareness, often resulting in biased outputs. However, comprehensive multilingual evaluation remains challenging due to limited benchmarks and questionable translation quality. To better assess these disparities, we introduce MakiEval, an automatic multilingual framework for evaluating cultural awareness in LLMs across languages, regions, and topics. MakiEval evaluates open-ended text generation, capturing how models express culturally grounded knowledge in natural language. Leveraging Wikidata’s multilingual structure as a cross-lingual anchor, it automatically identifies cultural entities in model outputs and links them to structured knowledge, enabling scalable, language-agnostic evaluation without manual annotation or translation. We then introduce four metrics that capture complementary dimensions of cultural awareness: granularity, diversity, cultural specificity, and consensus across languages. We assess 7 LLMs developed from different parts of the world, encompassing both open-source and proprietary systems, across 13 languages, 19 countries and regions, and 6 culturally salient topics (e.g., food, clothing). Notably, we find that models tend to exhibit stronger cultural awareness in English, suggesting that English prompts more effectively activate culturally grounded knowledge. We publicly release our code and data. [^1] https://github.com/mainlp/MAKIEval
@@ 1 Introduction
$\mathrm{ii}$) language and cultural coverage: most existing studies focus on English, adopt monolingual setups Onohara et al. (2025); Seveso et al. (2025); Naous et al. (2024); Nikandrou et al. (2025), or depend on translation-based pipelines for multilingual evaluation Berger and Ponti (2025). Although translation offers convenience, it introduces errors or meaning loss—particularly with culturally specific terms. Yao et al. (2024).
@@
To address these limitations, we introduce MakiEval, an automatic approach for $\mathrm{i}$) open-ended text evaluation, $\mathrm{ii}$) leveraging Wikidata as a language bridge and $\mathrm{iii}$) enabling structured evaluation metrics and scalable, up-to-date evaluation. Instead of static benchmarking, MakiEval is inspired from a human-centric perspective: when humans describe daily life in their own region, they tend to be more detailed (high granularity), mention a wider range of items (high diversity), refer to culturally specific concepts (high culture specificity), and maintain consistent representations across languages (high cultural consensus). These four metrics together capture key dimensions of cultural awareness in multilingual generation.
Through comprehensive experiments, we investigate how model origin, prompt language, and the mentioned country affect LLMs’ cultural awareness. We found models align more with culturally related languages and their regions of origin. Our contributions are:
@@ 2 Related Work
Some investigations examine abstract or high-level aspects of culture, such as value systems Ma et al. (2024); Agarwal et al. (2024); Shi et al. (2024), political ideology Bang et al. (2024), gender norms Wan et al. (2023), or expressions of stereotypes or hate Cheng et al. (2023); Deshpande et al. (2023). Others focus on concrete, content-level evaluations, assessing how LLMs engage with specific cultural knowledge or practices Naous et al. (2024); Wang et al. (2024); Jiang and Joshi (2024); Zhou et al. (2025); Hu et al. (2024). In this work, we also focus on the ideational branch, particularly on concrete cultural artifacts such as food, music, and clothing, which allow for a more structured, multilingual evaluation.
@@
To systematically evaluate cultural bias and awareness in LLMs, researchers have proposed a range of benchmarks by pre-collecting data and constructing dedicated benchmarks Yüksel et al. (2024); Jin et al. (2024); Chiu et al. (2024). Here, we highlight several recent works. JMMMU Onohara et al. (2025) focuses on Japanese culture awareness. ITALIC Seveso et al. (2025) targets cultural and linguistic understanding in Italian through a large-scale multiple-choice benchmark. CROPE Nikandrou et al. (2025) constructs a culturally grounded multimodal benchmark by collecting concepts from Wikipedia and generating 1,060 binary-choice English questions that ask whether an image corresponds to a culturally associated entity. Naous et al. (2024) analyze cultural bias in Arabic prompts by examining how language models balance Arab and Western representations across three tasks: stereotype generation (via adjective sentiment), fairness in NER and sentiment classification, and cultural entity prediction in masked tweets. Predicted entities are compared with the CAMeL Cultural Entities dataset, where overuse of Western entities signals bias.
@@ 3.2 Cultural Entity Extraction
We extract topic-relevant cultural entities from each generated text using GPT-4o-mini OpenAI (2025). For each topic (e.g., food, book or music), using GPT-4o-mini as the extraction model. It is then prompted to identify mentions of relevant entities in the text, i.e., dish names for the food topic or author names for the book topic. In addition to extraction, the model classifies each entity by its semantic granularity. For example, for the food topic, it distinguishes whether the mention refers to a specific dish (e.g., Kung Pao Chicken), a dish type (e.g., stir-fry), an individual ingredient (e.g., onion), or an ingredient class (e.g., vegetable).
@@ 3.3 Wikidata-Based Entity Match
Since Wikidata contains ambiguous labels (e.g., “lemon” may refer to a fruit or a film, each with a different QID), we implement a disambiguation procedure that leverages hierarchical semantic properties. For each matched label, we collect all candidate QIDs and perform a recursive traversal over their instance of, subclass of, and part of relations. Each candidate may be linked to multiple such properties, resulting in a tree of semantic paths. We traverse all reachable paths upward in the ontology graph and retain the entity if any path leads to a culturally relevant category (e.g. food, literary work). If no valid semantic path is found, the disambiguation is marked as unresolved.
@@
All matched entities and their corresponding metadata are compiled into a structured table that supports both quantitative and qualitative analysis. Each entry includes the entity’s QID, surface labels in multiple languages, granularity tags, and country/region-level information. We provide an evaluation of the performance in Appendix C.
@@ 4.1 Granularity
This metric captures the model’s ability to reference culturally specific concepts.
We define granularity as the level of cultural detail reflected in the predicted entities. Inspired by prior work on semantic abstraction in human descriptions Berger and Ponti (2025), we assign a granularity score to each entity based on a predefined cultural scale (e.g. music genre vs. song title, see Appendix B.2), with higher values indicating finer granularity. Let $\text{Gran}(e)$ be the granularity level of entity $e$. We compute:
@@ 4.2 Diversity
Diversity measures the breadth of cultural coverage by counting the number of unique cultural entities generated. As shown in Figure 2, the prompt asks about books, and the metric calculates the number of distinct book-related entities mentioned. Although the English response contains three extracted entities, one is repeated, so the resulting diversity score is 2. Higher diversity indicates that the model is capable of expressing a wide range of cultural knowledge for a given language and topic.
@@ 4.3 Culture Specificity
Culture specificity evaluates whether the entities in the generated text are aligned with the cultural context specified in the prompt. Figure 2 illustrates a case where the prompt inquires about the US. The culture specificity score is calculated by computing the proportion of entities associated with the United States among all extracted entities.
This metric reflects the model’s ability to generate culturally appropriate and contextually aligned content.
@@ 4.4 Culture Consensus
Culture consensus quantifies the agreement between model outputs for the same topic and cultural context setting (incl. same cultural context) across different languages. Let $Q_{A}$ and $Q_{B}$ denote the sets of Wikidata QIDs extracted from responses to the same prompt in languages $A$ and $B$, respectively. We compute their overlap using Jaccard similarity:
@@ 5.1 Models
Table 1: English prompt templates for the food category under three conditions: Neutral, which makes no cultural reference; Explicit and Implicit include a hints for referred culture. Full prompt templates for all categories are provided in the Appendix B.1.
We evaluate our framework on six multilingual language models, chosen to represent a diverse range of capabilities, resource requirements, and geographic origins. Our selection includes three open-source models developed across different regions: Llama-3.1-8B-Instruct (US) Grattafiori et al. (2024), Mistral-7B-Instruct-v0.1 (France) Jiang et al. (2023), and Qwen2.5-7B-Instruct (China) Yang et al. (2024). To complement these, we also include two proprietary large-scale models: ChatGPT-4o-mini (US) and DeepSeek-V3 (China) Liu et al. (2024a), which are widely used but not openly available or easily deployable on local infrastructure. In addition, we evaluate Llama-3.3-70B-Instruct Grattafiori et al. (2024) and aya-expanse-8b (Canada) Üstün et al. (2024). The inclusion of both Llama3 variants enables a controlled comparison of model size, while Aya allows us to assess cross-lingual generalization in a model explicitly trained for multilinguality.
@@ 5.2 Languages
We aim at a balanced and diverse coverage of major cultural regions and continents.
@@
Therefore, we select 13 languages as the primary focus of our study: Arabic, English, German, Hindi, Italian, Japanese, Korean, Persian, Simplified Chinese, Spanish, Thai, Traditional Chinese, and Turkish. The chosen languages span multiple language families (e.g., Indo-European, Sino-Tibetan), writing systems (e.g., Kanji, Devanagari), and geographic regions (for example, Middle East). [^2] In addition, we initially explored two low-resource African languages: Hausa and Yoruba. However, during experimentation, we found that open-source models produced text of very poor quality in these two languages, often entirely unrelated to the prompts (see Appendix B). As a result, we set aside experiments involving native African languages for this work.
@@ 5.3 Prompts
We design three categories of prompts to elicit culturally grounded responses from language models: explicit contextualized prompts, where the cultural background is explicitly (e.g. “about a person from Spain”), implicit contextualized prompts (e.g. “about Juan”, where “Juan” is a popular name in Spain) and the neutral prompts with no mention of nationality, ethnicity, or location, allowing us to assess cultural preferences expressed by the model without direct cues. Table 1 provides examples of contextualized and neutral prompt templates.
@@ 6.1 Granularity
Differences do exist when moving with the same model to a different prompt language. E.g., English and Turkish have lower granularity levels for most models compared to Chinese. The observed patterns may stem from inherent linguistic differences in lexical specificity when describing cultural concepts, where Chinese, e.g., might employ more fine-grained expressions. Or alternatively, these differences might reflect limitations in the models’ multilingual capabilities.
@@ 6.2 Diversity
Figure 5: Diversity metric comparing prompts with culture context India in the native language Hindi vs. other languages. Each row shows a separate cultural topic.
The diversity metric varies widely depending on the setting, ranging from a handful of distinct entities per topic to over a hundred. The differences depend on both the prompt language and the model with Llama3 having the highest mean diversity of 36.65 and DeepSeek the lowest with 14.62. In English for the book topic and the United States, we even observe a diversity of nearly 1, outputting the entity To Kill a Mockingbird in 499 out of 500 generations.
@@ 6.3 Cultural Specificity
Prompting with the explicit cultural context increases the cultural specificity in most cases (Figure 4 right). Implicit prompts (that do not mention the country explicitly) in LLaMA3 produce weaker gains than explicit ones (Figure 21 in Appendix), indicating reliance on overt cultural cues. Crucially, English prompts consistently achieve the highest specificity gains across nearly all countries and models (see Appendix F). This suggests that LLMs are more sensitive to cultural signals in English, likely due to the disproportionate amount of English data in their pretraining corpora. This further underscores the importance of multilingual evaluation, as cultural awareness is not only model-dependent but also deeply shaped by language.
@@ 7 Conclusion
Our results highlight that cultural awareness in LLMs is more complex than simple descriptions like a general bias to Western culture. Rather, it manifests in different forms, like the granularity, specificity and diversity of cultural entities. Besides general model capabilities, we show that the prompting language plays a crucial role in evoking these cultural differences, even when no explicit cultural context is given. The sensitivity to cultural signals in English, the connection between countries and their local languages, and the cultural consensus that link regional areas emphasize the need for multilingual and multicontextual evaluation of cultural awareness.
To foster further analysis of the cultural awareness of LLMs, we will release the full dataset of generated outputs and cultural entities for all model, language, country and topic combinations. Our work also provides guidance for future research in identifying how these factors have been picked up by the models during training and alignment, and how culture is represented internally in the models.
Acknowledgments
We thank the members of MaiNLP for their valuable feedback on this project, especially Yupei Du, Bolei Ma, Verena Blaschke, Elena Senger, Florian Echin, Soh-Eun Shim and Rob van der Goot.
@@ Acknowledgments
First, our experiments are limited to a subset of high- and mid-resource languages. Although Wikidata provides labels in many low-resource languages, we exclude them due to the low reliability of generation from current language models in these settings, which makes meaningful evaluation infeasible.
@@
We do not foresee any ethical concerns associated with this work. All analyses were conducted using publicly available models. No private or sensitive information was used. The texts we studied were also model-generated and do not represent our views. Additionally, we will release our code, produced data and documentations to support transparency and reproducibility.

### v3

# Raoyuan Zhao, Beiduo Chen, Barbara Plank, Michael A. Hedderich

@@ Abstract
Large language models (LLMs) are used globally across many languages, but their English-centric pretraining raises concerns about cross-lingual disparities for cultural awareness, often resulting in biased outputs. However, comprehensive multilingual evaluation remains challenging due to limited benchmarks and questionable translation quality. To better assess these disparities, we introduce MakiEval, an automatic multilingual framework for evaluating cultural awareness in LLMs across languages, regions, and topics. MakiEval evaluates open-ended text generation, capturing how models express culturally grounded knowledge in natural language. Leveraging Wikidata’s multilingual structure as a cross-lingual anchor, it automatically identifies cultural entities in model outputs and links them to structured knowledge, enabling scalable, language-agnostic evaluation without manual annotation or translation. We then introduce four metrics that capture complementary dimensions of cultural awareness: granularity, diversity, cultural specificity, and consensus across languages. We assess 7 LLMs developed from different parts of the world, encompassing both open-source and proprietary systems, across 13 languages, 19 countries and regions, and 6 culturally salient topics (e.g., food, clothing). Notably, we find that models tend to exhibit stronger cultural awareness in English, suggesting that English prompts more effectively activate culturally grounded knowledge.
@@ 1 Introduction
$\mathrm{ii}$) language and cultural coverage: most existing studies focus on English, adopt monolingual setups Onohara et al. (2025); Seveso et al. (2025); Naous et al. (2024); Nikandrou et al. (2025), or depend on translation-based pipelines for multilingual evaluation Berger and Ponti (2025). Although translation offers convenience, it introduces errors or meaning loss—particularly with culturally specific terms Yao et al. (2024).
@@
To address these limitations, we introduce MakiEval, an automatic approach for $\mathrm{i}$) open-ended text evaluation, $\mathrm{ii}$) leveraging Wikidata as a language bridge and $\mathrm{iii}$) enabling structured evaluation metrics and scalable, up-to-date evaluation. Instead of static benchmarking, MakiEval is inspired from a human perspective and prior work: when humans describe daily life in their own region, they tend to be more detailed (high granularity), mention a wider range of items (high diversity), refer to culturally specific concepts (high culture specificity), and maintain consistent representations across languages (high cultural consensus) Chafe (1994); Dewaele and Pavlenko (2003); Peters et al. (2018). For example, when asked about Chinese breakfast, an English speaker unfamiliar with the culture might respond simply “A Chinese eats noodles in the morning,” whereas a Chinese speaker would provide a richer description such as “youtiao (fried dough sticks) with soy milk, jianbing, or millet porridge in winter.” This contrast highlights all four dimensions. These dimensions have long been discussed in psycholinguistics and linguistics, and they also resonate with recent NLP research on semantic abstraction, coverage, cultural awareness, and cross-lingual consistency. In what follows, we draw on this line of work to design four operational metrics that capture cultural awareness in multilingual generation.
Building on these metrics, we conduct comprehensive experiments to investigate how model origin, prompt language, and the mentioned country affect LLMs’ cultural awareness. We found models align more with culturally related languages and their regions of origin. We publicly release our code and data. [^1] https://github.com/mainlp/MAKIEval Our contributions are:
@@ 2 Related Work
Some investigations examine abstract or high-level aspects of culture, such as value systems Ma et al. (2024); Agarwal et al. (2024); Shi et al. (2024), political ideology Bang et al. (2024), gender norms Wan et al. (2023), or expressions of stereotypes or hate Cheng et al. (2023); Deshpande et al. (2023). Others focus on concrete, content-level evaluations, assessing how LLMs engage with specific cultural knowledge or practices Naous et al. (2024); Wang et al. (2024); Jiang and Joshi (2024); Zhou et al. (2025); Hu et al. (2024); Ghaboura et al. (2025). In this work, we also focus on the ideational branch, particularly on concrete cultural artifacts such as food, music, and clothing, which allow for a more structured, multilingual evaluation.
@@
To systematically evaluate cultural bias and awareness in LLMs, researchers have proposed a range of benchmarks by pre-collecting data and constructing dedicated benchmarks Yüksel et al. (2024); Jin et al. (2024); Chiu et al. (2024); Singh et al. (2025). Here, we highlight several recent works. JMMMU Onohara et al. (2025) focuses on Japanese culture awareness. ITALIC Seveso et al. (2025) targets cultural and linguistic understanding in Italian through a large-scale multiple-choice benchmark. CROPE Nikandrou et al. (2025) constructs a culturally grounded multimodal benchmark by collecting concepts from Wikipedia and generating 1,060 binary-choice English questions that ask whether an image corresponds to a culturally associated entity. Naous et al. (2024) analyze cultural bias in Arabic prompts by examining how language models balance Arab and Western representations across three tasks: stereotype generation (via adjective sentiment), fairness in NER and sentiment classification, and cultural entity prediction in masked tweets. Predicted entities are compared with the CAMeL Cultural Entities dataset, where overuse of Western entities signals bias.
@@ 3.2 Cultural Entity Extraction
We extract topic-relevant cultural entities from each generated text using GPT-4o-mini OpenAI (2025). For each topic (e.g., food, books, or music), GPT-4o-mini is prompted to identify mentions of relevant entities in the text (e.g., dish names for the food topic or author names for the book topic). In addition to extraction, the model classifies each entity by its semantic granularity. For example, for the food topic, it distinguishes whether the mention refers to a specific dish (e.g., Kung Pao Chicken), a dish type (e.g., stir-fry), an individual ingredient (e.g., onion), or an ingredient class (e.g., vegetable).
@@ 3.3 Wikidata-Based Entity Match
Since Wikidata contains ambiguous labels (e.g., “lemon” may refer to a fruit or a film, each with a different QID), we implement a disambiguation procedure that leverages hierarchical semantic properties. For each matched label, we collect all candidate QIDs and perform a recursive traversal over their instance of, subclass of, and part of relations. Each candidate may be linked to multiple such properties, resulting in a tree of semantic paths. We traverse all reachable paths upward in the ontology graph and retain the entity if any path leads to a culturally relevant category (e.g. food, literary work). If no valid semantic path is found, the disambiguation is marked as unresolved, and the corresponding entity is discarded from evaluation.
@@
All matched entities and their corresponding metadata are compiled into a structured table that supports both quantitative and qualitative analysis. Each entry includes the entity’s QID, surface labels in multiple languages, granularity tags, and country/region-level information. A manual evaluation of the performance indicated strong results, with full details presented in Appendix C.
@@ 4.1 Granularity
This metric captures the model’s ability to reference culturally specific concepts. We define granularity as the level of cultural detail reflected in the predicted entities. Inspired by a prior work on semantic abstraction in human descriptions Berger and Ponti (2025), we assign a granularity score to each entity based on a predefined cultural scale (e.g. music genre vs. song title, see Appendix B.2), with higher values indicating finer granularity. Let $\text{Gran}(e)$ be the granularity level of entity $e$. We compute:
@@ 4.2 Diversity
Diversity measures the breadth of cultural coverage by counting the number of unique cultural entities generated and is inspired by and adapted from work by Berger and Ponti (2025). As shown in Figure 2, the prompt asks about books, and the metric calculates the number of distinct book-related entities mentioned. Although the English response contains three extracted entities, one is repeated, so the resulting diversity score is 2. Higher diversity indicates that the model is capable of expressing a wider range of cultural knowledge for a given language and topic.
@@ 4.3 Culture Specificity
Culture specificity evaluates whether the entities in the generated text are aligned with the cultural context specified in the prompt. Figure 2 illustrates a case where the prompt inquires about the US. The culture specificity score is calculated by computing the proportion of entities associated with the United States among all extracted entities. This metric reflects the model’s ability to generate culturally appropriate and contextually aligned content. Its design is motivated by earlier studies on cultural awareness and bias Wang et al. (2024); Liu et al. (2024b).
@@ 4.4 Culture Consensus
This metric is motivated by the observations of Hedderich et al. (2025), who showed that LLMs prompted with the same query in different languages often yield inconsistent outputs, highlighting the need to evaluate whether models preserve shared cultural understanding across languages. Culture consensus quantifies the agreement between model outputs for the same topic and cultural context setting across different languages. Let $Q_{A}$ and $Q_{B}$ denote the sets of Wikidata QIDs extracted from responses to the same prompt in languages $A$ and $B$, respectively. We compute their overlap using Jaccard similarity:
@@ 5.1 Models
Table 1: English prompt templates for the food category under three conditions: Neutral, which makes no cultural reference; Explicit and Implicit include a hint for the referred cultural context. Full prompt templates for all categories are provided in Appendix B.1.
We evaluate our framework on seven multilingual language models, chosen to represent a diverse range of capabilities, resource requirements, and geographic origins. Our selection includes three open-source models developed across different regions: Llama-3.1-8B-Instruct (US) Grattafiori et al. (2024), Mistral-7B-Instruct-v0.1 (France) Jiang et al. (2023), and Qwen2.5-7B-Instruct (China) Yang et al. (2024). To complement these, we also include two proprietary large-scale models: ChatGPT-4o-mini (US) and DeepSeek-V3 (China) Liu et al. (2024a), which are widely used but not openly available or easily deployable on local infrastructure. In addition, we evaluate Llama-3.3-70B-Instruct Grattafiori et al. (2024) and aya-expanse-8b (Canada) Üstün et al. (2024). The inclusion of both Llama3 variants enables a controlled comparison of model size, while Aya allows us to assess cross-lingual generalization in a model explicitly trained for multilinguality.
@@ 5.2 Languages
We aim at a balanced and diverse coverage of major cultural regions and continents. Therefore, we select 13 languages as the primary focus of our study: Arabic, English, German, Hindi, Italian, Japanese, Korean, Persian, Simplified Chinese, Spanish, Thai, Traditional Chinese, and Turkish. The chosen languages span multiple language families (e.g., Indo-European, Sino-Tibetan), writing systems (e.g., Kanji, Devanagari), and geographic regions (for example, Middle East). Table 2 provides an overview of the selected languages and their corresponding regions [^2] In addition, we initially explored two low-resource African languages: Hausa and Yoruba. However, during experimentation, we found that open-source models produced text of very poor quality in these two languages, often entirely unrelated to the prompts (see Appendix B). As a result, we set aside experiments involving native African languages as future work.
@@ 5.3 Prompts
We design three categories of prompts to elicit culturally grounded responses from language models: explicit contextualized prompts, where the cultural background is explicitly mentioned (e.g. “about a person from Spain”), implicit contextualized prompts, where a culturally associated name is mentioned (e.g. “about Juan”, where “Juan” is a popular name in Spain) and the neutral prompts with no mention of nationality, ethnicity, or location, allowing us to assess cultural preferences expressed by the model without direct cues. Table 1 provides examples of contextualized and neutral prompt templates.
@@ 6 Results and Analysis
Figure 4: Diversity metric comparing prompts with culture context India in the native language Hindi vs. other languages for Llama3. Each row shows a separate cultural topic.
@@ 6.1 Granularity
Differences exist when moving with the same model to a different prompt language. E.g., English and Turkish have lower granularity levels for most models compared to Chinese. The observed patterns may stem from inherent linguistic differences in lexical specificity when describing cultural concepts, where Chinese, e.g., might employ more fine-grained expressions. Or alternatively, these differences might reflect limitations in the models’ multilingual capabilities.
@@ 6.2 Diversity
The diversity metric varies widely depending on the setting, ranging from a handful of distinct entities per topic to over a hundred. The differences depend on both the prompt language and the model, with Llama3 having the highest mean diversity of 36.65 and DeepSeek the lowest with 14.62. For DeepSeek, when prompted in English on the book topic for the United States, we observe a diversity score close to 1, with the model producing the entity To Kill a Mockingbird in 499 out of 500 generations.
@@ 6.3 Cultural Specificity
Prompting with the explicit cultural context increases the cultural specificity in most cases (Figure 5 right). Implicit prompts (that do not mention the country explicitly) in LLaMA3 produce weaker gains than explicit ones (Figure 21 in Appendix), indicating reliance on overt cultural cues. Crucially, English prompts consistently achieve the highest specificity gains when adding cultural context across nearly all countries and models (see Appendix F). This suggests that LLMs are more sensitive to cultural signals in English, likely due to the disproportionate amount of English data in their pretraining corpora. This further underscores the importance of multilingual evaluation, as cultural awareness is not only model-dependent but also deeply shaped by language.
@@ 6.4 Cultural Consensus
7 Discussion
The design of the metrics in MakiEval is informed by insights from psycholinguistics and linguistics, as well as prior work in NLP, and our experiments demonstrate systematic differences across languages and models. However, how these differences should be interpreted and applied in practice remains an open question that merits further discussion. MakiEval is not intended as a prescriptive benchmark in which higher values are inherently preferable across all dimensions. Instead, it provides a set of complementary measures to capture different aspects of how LLMs represent and express cultural knowledge.
The relevance and desirability of these measures vary with the characteristics of the downstream task and the interpretation of the proposed metrics requires careful consideration. For example, high specificity and granularity may be beneficial in settings where interlocutors share a cultural background, while lower values may be more effective in intercultural communication where common ground cannot be assumed. Similarly, cross-lingual consensus may indicate robustness in knowledge-intensive applications such as multilingual question answering and cross-lingual understanding Chen et al. (2022b); Qi et al. (2023); Wang et al. (2025), whereas divergence can be advantageous in tasks that benefit from creativity or cultural variation, including storytelling or adaptation to local contexts. Given the diversity of generation tasks, no single value or direction of change should be considered universally optimal. Instead, the task-specific needs should be considered along the the proposed cultural dimensions.
@@ 8 Conclusion
Our results highlight that cultural awareness in LLMs is more complex than simple descriptions like a general bias to Western culture. Rather, it manifests in different forms, like the granularity, specificity and diversity of cultural entities. Besides general model capabilities, we show that the prompting language plays a crucial role in evoking these cultural differences, even when no explicit cultural context is given. The sensitivity to cultural signals in English, the connection between countries and their local languages, and the cultural consensus that links regional areas emphasize the need for multilingual and multicontextual evaluation of cultural awareness.
To foster further analysis of the cultural awareness of LLMs, we will release the full dataset of generated outputs and cultural entities for all model, language, country and topic combinations. Our work also provides guidance for future research in identifying how these factors have been picked up by the models during training and alignment, and how culture is represented internally in the models. More broadly, our framework offers a generalizable paradigm for multilingual evaluation, our findings reveal systematic disparities that call for rethinking prompting and training beyond English, and our metrics provide actionable signals for improving fairness and inclusivity in global LLM applications. We hope this work contributes to more equitable evaluation and deployment of language technologies worldwide.
@@
First, our experiments are limited to a subset of high- and mid-resource languages. Although Wikidata provides labels in many low-resource languages, we excluded low-resource languages due to the low generation quality from current language models in these settings.
@@
We do not foresee any ethical concerns associated with this work. All analyses were conducted using publicly available models. No private or sensitive information was used. The texts we studied were model-generated and do not necessarily represent our views. Additionally, we will release our code, produced data and documentations to support transparency and reproducibility.
Acknowledgments
We thank the members of MaiNLP for their valuable feedback on this project, especially Verena Blaschke, Yupei Du, Florian Echin, Rob van der Goot, Bolei Ma, Elena Senger, and Soh-Eun Shim. We are further grateful to a number of these colleagues, as well as to Fatima Abdulla Mohammed Obaid Almatrooshi, Tianshi Feng, Paravee Jungbauer, Neysa Kaustubh, Abdullatif Köksal, Yuchen Li, Ali Modarressi, and Jing’an Qian for their assistance in verifying the language-specific texts studied in this paper.